OpenStudy (anonymous):

Simplify 3/(x-7)-1/(x^2-49)+3/(x^2-14x+49) a.x^2-x-24/(x-7)(x+7)^2 b.3x^2+x+24/(x-7)(x+7)^2 c.3x^2+2x-119/(x-7)^2(x+7) d.3x^2-2x-49/(x-7)^2(x+7)

OpenStudy (lalaly):

\[\frac{3}{(x-7)}-\frac{1}{x^2-49}+\frac{3}{x^2-14x+49}\]\[=\frac{3}{x-7}-\frac{1}{(x-7)(x+7)}+\frac{3}{(x-7)^2}\]now find a common denominator\[(\frac{3}{(x-7)} \times \frac{(x-7)(x+7)}{(x-7)(x+7)}) - (\frac{1}{(x-7)(x+7)} \times \frac{x-7}{x-7}) +(\frac{3}{(x-7)^2} \times \frac{x+7}{x+7})\]\[\frac{3(x-7)(x+7)-(x-7)+3(x+7)}{(x-7)^2(x+7)}\]

OpenStudy (anonymous):

lalaly how do you create a fraction in this equation thingy.

OpenStudy (lalaly):

u can use the equation button down there ... or u can write in latex right away

OpenStudy (anonymous):

lalaly that answer is not found in one of the answers on my test

OpenStudy (lalaly):

lol it needs more simplification wait

OpenStudy (anonymous):

time to learn latex~

OpenStudy (anonymous):

ok